[The following information applies to the questions displayed below.] Oslo Company prepared the following contribution format6. If the selling price increases by $2 per unit and the sales volume decreases by 100 units, what would be the net operating7. If the variable cost per unit increases by $1, spending on advertising increases by $1,500, and unit sales increase by 2008. What is the break-even point in unit sales? Break-even point units9. What is the break-even point in dollar sales? Break-even point10. How many units must be sold to achieve a target profit of $12,600? Number of units
Answer #1

6) Net operating income

Total Per unit
Sales (62*900) 55800 62
Variable cost 35100 39
Contribution margin 20700 23
Fixed cost 14700
Net operating income 6000

7) Net operating income

Total Per unit
Sales (1200*60) 72000 60
Variable cost 48000 40
Contribution margin 24000 20
Fixed cost 16200
Net operating income 7800

8) Break even unit = 14700/21 = 700 Units

9) Break even sales = 700*60 = $42000

10)Required unit = (14700+12600)/21 = 1300 Units